1. It is given that three moles of helium gas are at a temperature of 32 °C and at a pressure of 15 atm. Helium gas is heated while being at a constant volume container. As a result, the pressure of gas triples. what is the final temperature of it


1
Expert's answer
2021-05-19T20:46:08-0400

as V , n , R are constant

in PV=nRT

"P_1=15 \\
T_1=32 ^o C=305 K \\ \ \ \ \ \ ^oC \ to \ K(kelvin)\ by \ adding \ 273 \\
P_2=3\times15=45\ atm"


so we will use the relation

"\frac{P_1}{T_1}=\frac{P_2}{T_2}"

"T_2=P_2\frac{T_1}{P_1}"

"T_2=45\times\frac{305}{15}=915K=642^oC"